1. Something is wrong with the student’s work below.
• Explain what the student did wrong.
• Next, show how to simplify the fraction correctly.
Sara simplifies the expression as follows: x^2+6x+9/x+3
x^2 + 6x( X is crossed out with 1)+9( 9 is crossed out with 3/ x (crossed out for 1)+3( 3 crossed out for 1)=x^2+9( x does not eqaul -3)


Answer by richwmiller(17219)

(x^2+6x+9)=(x+3)*(x+3)
after the division we have x+3
(x+3)*(x+3)/(x+3)=x+3